the 100000 loan has to be repaid after 21 years. determine the amount of the 21st and final payment when 20 payments of 10,000 have been made in each year. the interest rate remains at 8 % per year?
Answer #1

Loan amount=$100000

Annual installment=$10000 for 20 years

Rate of interest=i=8%

Balance left after 20 installments=100000*(F/P,0.08,20)-10000*(F/A,0.08,20)

Let us calculate the interest factors

(F/A,0.08,20)=(1+0.08)20=4.66095714

Balance left after 20 installments=100000*4.66095714-10000*45.76196430=8476.07

Payment to be made at the end of 21st year=8476.07*(1+8%)=$9154.16